OpenStudy (anonymous):

There are enough oranges in a basket to give 12 oranges to each of 20 children, with no oranges left over. If 5 children do not want any oranges, how many oranges can be given to each of the others?

OpenStudy (anonymous):

so there are how many total oranges-- 12*20=240

OpenStudy (savvy):

12*20/15

OpenStudy (anonymous):

now 240 / 15=16

OpenStudy (savvy):

sicnce total oranges = 12*20 to each of 15 students, = total/15

OpenStudy (anonymous):

true